# Nonclassical dynamics of Néel vector and magnetization accompanied by THz and high-harmonic radiation from ultrafast-light-driven NiO antiferromagnet insulator

### v3

@@ Abstract
Introduction.—Pump-probe experiments [1] on strongly correlated antiferromagnetic (AF) insulators (AFI), like prototypical NiO, have revealed [2] intricate fundamental effects, such as coherent long-lived many-body dynamics resulting from a coupling to the AF background where coherent oscillations persists on surprisingly long timescales ps. The fs laser pulse (fsLP) in these experiments has a central frequency that is above-gap between two Hubbard bands [Fig. 2], as it has also been typically the focus of theoretical studies [2, 3]. Theoretical interest also exists to understand quantum tunneling, multiphoton absorption and the so-called “Keldysh crossover” and the ensuing nonlinear doublon-holon pair production in the case of subgap fsLP pump [4, 5, 6, 7]. These findings comprise a panoply of fundamental effects that cannot be found in fsLP-driven conventional band insulators and semiconductors.
@@
The same NiO material driven by subgap fsLP, but in combination with heavy metal (HM) layer (like Pt,W,Ta) introducing effects [8] due to strong spin-orbit coupling (SOC), have been very recently explored [9, 10] as spintronic THz emitter [11, 12, 13, 14, 15, 16]. Isolated ferromagnetic metal (FM) layers (such as Co,Fe,Ni) or FM/HM bilayers have been intensely studied for nearly 30 years in order to understand ultrafast demagnetization [17, 18], as well as THz emission by these systems, which is much stronger in the case of bilayers [11, 12, 13, 14, 15, 19] than in the case of a single FM layer [20, 14, 19]. The fsLP in spintronic experiments typically has a central wavelength of nm, so that its photons have energy centered around eV which is subgap [9, 10, 21] with respect to the separation eV between the two Hubbard bands of NiO. The spintronic phenomena in such experiments on subgap-fsLP-excited NiO have been interpreted [9, 10] by borrowing the standard intuitive picture [11, 12, 13, 14, 15, 19] (for its recent modifications, however, via microscopic theory see Refs. [22, 23, 24]) developed for FM/HM bilayers. That is, an ultrafast spin current is somehow generated that flows from NiO into the HM layer, so that latter can convert it into charge current via the inverse spin Hall effect [25]. The time-dependent charge current is needed to obtain sizable electromagnetic (EM) radiation in the far-field (FF) region [Fig. 1], as well as to interpret [9, 10] the enhancement [14] of the emitted radiation when switching from FM to FM/HM systems. This is because magnetic dipole radiation [20, 14, 19] from time-dependent magnetization is smaller [22] than radiation by a time-dependent charge current. However, this picture does not explain the microscopic mechanism that generates the assumed spin current in NiO/HM bilayers (only speculation exist thus far [26]), or why the frequency spectrum of such current has features within 0.1–3 THz range that is imprinted in the EM radiation detected experimentally [9, 10]. It is obvious that fsLP will drive electrons into dynamics at its own center frequency , as well as at integer (typically odd [27, 28]) multiples of for sufficiently intense fsLP. Such high-harmonic generation (HHG) in current and EM radiation has been vigorously explored in recent years in diverse quantum materials [29, 30], including strongly correlated ones [6, 7, 31, 32, 27]. Finally, it remains unclear what type of dynamics is obeyed by the Néel vector (as the difference of two sublattices magnetizations) and magnetization [as the sum of sublattice magnetizations, which is necessarily a nonequilibrium quantity because ], when compared to standard demagnetization [17, 18, 14] of FM layers, where magnetization vector shrinks. In the case of thin FM layers, a rapid and straightforward analysis of the direction of and its magnitude is achieved via [14] magneto-optical Kerr or Faraday effects. In contrast, they do not apply to AFs, so novel ideas have been explored [33] to detect the presumed [34, 35, 10] rotation of the Néel vector.
@@
Thus, developing a microscopic understanding of the response of AFI to subgap fsLP (that is, by starting from a suitable quantum many-body Hamiltonian and using tools of nonequilibrium quantum statistical mechanics), would also help to resolve a number of outstanding issues in AF optospintronics [42]. Note that, specifically for NiO which is a strongly correlated material sharing features of both Mott and charge-transfer insulators [2, 43], X-ray techniques applied as a probe after subgap fsLP pump have revealed [21] possible substantial [1, 3] changes of its electronic structure, such as emergence of midgap states and Hubbard gap reduction persisting on time scales ps. Such phenomena originating from charge dynamics must be taken into account together with local spin dynamics, as they can lead to inextricable complex spin-charge dynamics [2, 44]. In the case of weakly correlated FMs, proper description of spin-charge dynamics is achieved via time-dependent density functional theory (TDDFT) which has provided [45, 46, 47, 48, 49] a most detailed insight into a sequence of fast changing events [50] and their effect [22, 23] on THz emission. However, application [51] of TDDFT to NiO is impeded by intricacies in including a time-dependent [52, 53] Hubbard to properly capture a strong on-site Coulomb interaction that is also dynamical (as opposed to static in conventional equilibrium DFT+U [54] calculations).
@@
By considering a multi-orbital ladder, spin-charge separation is largely suppressed, so a plethora of dissipative mechanisms based on electron-spin interaction become active, such as the AF background interaction with doublons and holons [2, 70] or local dissipation caused by the Hund interaction [71, 72, 73]. Our principal results, revealing highly nonclassical dynamics of Néel vector and nonequilibrium magnetization, and the ensuing radiation at both high harmonics of fsLP and in much lower THz range, are given in Figs. 3 and 4. Prior to discussing these results, we introduce useful concepts and notations.

### v4

@@ Abstract
Introduction.—Pump-probe experiments [1] with strongly correlated antiferromagnetic (AF) insulators (AFI), like prototypical NiO [2], reveal [3] exotic effects interweaving nonequilibrium many-body physics and quantum coherence that can persist on surprisingly long timescales (such as ps [2]) due to a large Mott gap [Fig. 2] providing protection against fast thermalization and heating. The fs laser pulse (fsLP) in these experiments and typical theoretical studies [2, 4] has a central frequency that is above-gap between two Hubbard bands [Fig. 2]. Theoretical interest also exists to understand quantum tunneling, multiphoton absorption and the so-called “Keldysh crossover” and the ensuing nonlinear doublon-holon pair production in the case of subgap fsLP pump [5, 6, 7, 8]. Such panoply of complex nonequilibrium many-body states [3] cannot be found in fsLP-driven conventional band insulators and semiconductors, where single particle [9] description suffices.
@@
The same NiO material driven by subgap fsLP, but in combination with heavy metal (HM) layer (like Pt,W,Ta) introducing effects [10] due to strong spin-orbit coupling (SOC), have been very recently explored [11, 12] as spintronic THz emitter [13, 14, 15, 16, 17, 18]. Isolated ferromagnetic metal (FM) layers (such as Co,Fe,Ni) or FM/HM bilayers have been intensely studied for nearly 30 years in order to understand ultrafast demagnetization [19, 20], as well as THz emission by these systems, which is much stronger in the case of bilayers [13, 14, 15, 16, 17, 21] than in the case of a single FM layer [22, 16, 21]. The fsLP in spintronic experiments typically has a central wavelength of nm, so that its photons have energy centered around eV which is subgap [11, 12, 23] with respect to the separation eV between the two Hubbard bands of NiO. The spintronic phenomena in such experiments on subgap-fsLP-excited NiO have been interpreted [11, 12] by borrowing the standard intuitive picture [13, 14, 15, 16, 17, 21] (for its recent modifications, however, via microscopic theory see Refs. [24, 25, 26]) developed for FM/HM bilayers. That is, an ultrafast spin current is somehow generated that flows from NiO into the HM layer, so that latter can convert it into charge current via the inverse spin Hall effect [27]. The time-dependent charge current is needed to obtain sizable electromagnetic (EM) radiation in the far-field (FF) region [Fig. 1], as well as to interpret [11, 12] the enhancement [16] of the emitted radiation when switching from FM to FM/HM systems. This is because magnetic dipole radiation [22, 16, 21] from time-dependent magnetization is smaller [24] than radiation by a time-dependent charge current. However, this picture does not explain the microscopic mechanism that generates the assumed spin current in NiO/HM bilayers (only speculation exist thus far [28]), or why the frequency spectrum of such current has features within 0.1–3 THz range that is imprinted in the EM radiation detected experimentally [11, 12]. It is obvious that fsLP will drive electrons into dynamics at its own center frequency , as well as at integer (typically odd [29, 30]) multiples of for sufficiently intense fsLP. Such high-harmonic generation (HHG) in current and EM radiation has been vigorously explored in recent years in diverse quantum materials [31, 32], including strongly correlated ones [7, 8, 33, 34, 29, 3]. Finally, it remains unclear what type of dynamics is obeyed by the Néel vector (as the difference of two sublattices magnetizations) and magnetization [as the sum of sublattice magnetizations, which is necessarily a nonequilibrium quantity because ], when compared to standard demagnetization [19, 20, 16] of FM layers, where magnetization vector shrinks. In the case of thin FM layers, a rapid and straightforward analysis of the direction of and its magnitude is achieved via [16] magneto-optical Kerr or Faraday effects. In contrast, they do not apply to AFs, so novel ideas have been explored [35] to detect the presumed [36, 37, 12] rotation of the Néel vector.
@@
Thus, developing a microscopic understanding of the response of AFI to subgap fsLP (that is, by starting from a suitable quantum many-body Hamiltonian and using tools of nonequilibrium quantum statistical mechanics), would also help to resolve a number of outstanding issues in AF optospintronics [44]. Note that, specifically for NiO which is a strongly correlated material sharing features of both Mott and charge-transfer insulators [2, 45], X-ray techniques applied as a probe after subgap fsLP pump have revealed [23] possible substantial [1, 4] changes of its electronic structure, such as emergence of midgap states and Hubbard gap reduction persisting on time scales ps. Such phenomena originating from charge dynamics must be taken into account together with local spin dynamics, as they can lead to inextricable complex spin-charge dynamics [2, 46]. In the case of weakly correlated FMs, proper description of spin-charge dynamics is achieved via time-dependent density functional theory (TDDFT) which has provided [47, 48, 49, 50, 51] a most detailed insight into a sequence of fast changing events [52] and their effect [24, 25] on THz emission. However, application [53] of TDDFT to NiO is impeded by intricacies in including a time-dependent [54, 55, 3] Hubbard to properly capture a strong on-site Coulomb interaction that is also dynamical (as opposed to static in conventional equilibrium DFT+U [56] calculations).
@@
By considering a multi-orbital ladder, spin-charge separation is largely suppressed, so a plethora of dissipative mechanisms based on electron-spin interaction become active, such as the AF background interaction with doublons and holons [2, 3] or local dissipation caused by the Hund interaction [72, 73, 74]. Our principal results, revealing highly nonclassical dynamics of Néel vector and nonequilibrium magnetization, and the ensuing radiation at both HHG of fsLP and in much lower THz range, are given in Figs. 3 and 4. Prior to discussing these results, we introduce useful concepts and notations.